Owen spent 1/3 of his monthly pay check at the movies. He also received a $25 gift certificate to the movies from his parents. If he spent a total of $45 at the movie theatre this month how much was his pay check ?
Answer by rfer(16322)

45-25=20
20/(1/3)=
20*(3/1)=$60